40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 25 19:42:12 2014, ericq, Update, General, Vent progress 
    Reply  Fri Sep 26 12:25:34 2014, ericq, Update, General, Vent progress 
       Reply  Fri Sep 26 16:10:14 2014, ericq, Update, General, Vent update 
          Reply  Fri Sep 26 17:13:39 2014, ericq, Update, General, Vent update 
             Reply  Mon Sep 29 12:47:51 2014, ericq, Update, General, Vent update 
                Reply  Mon Sep 29 17:10:51 2014, ericq, Update, General, Vent update 
                   Reply  Mon Sep 29 18:12:24 2014, ericq, Update, General, Vent update 
                      Reply  Tue Sep 30 11:53:29 2014, ericq, Update, General, Vent update 
                         Reply  Tue Sep 30 16:18:57 2014, Steve, Update, VAC, pump down #78 has started pd78hastarted.png
                            Reply  Tue Sep 30 18:02:53 2014, Steve, Update, VAC, pump down #78 stops at 320 Torr 3h.png
                               Reply  Wed Oct 1 10:21:19 2014, Steve, Update, VAC, pump down #78 stops at 3.5 Torr pd78secondStop.png
                                  Reply  Wed Oct 1 16:26:53 2014, Steve, Update, VAC, pump down #78 completed pd78completedin25hrs.png
                                     Reply  Thu Oct 2 09:23:23 2014, Steve, Update, VAC, vent 77 vent77SRMmoved.png
                                        Reply  Fri Oct 3 10:10:37 2014, Steve, Update, VAC, cold cathode gauge reading switched CC1switched.png
                                           Reply  Fri May 8 14:30:09 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now cc1bake95C.jpg
                                              Reply  Tue May 12 14:57:52 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now 
Message ID: 10544     Entry time: Fri Sep 26 12:25:34 2014     In reply to: 10541     Reply to this: 10545
Author: ericq 
Type: Update 
Category: General 
Subject: Vent progress 

I figured out that didn't change the initial guess for the fit routine in Gabriele's code. I also changed the fminsearch criteria to least squares fitting, instead of minimax. The consistency checks now look just as good as the previous time we did these kind of measurements, no disagreements bigger than 1.6mm. 

Thus, the current estimate of the SRC length after yesterday's motions is 5402mm, where we desire 5399mm. So, we will try to move SRM 3mm closer to SR2, after confirming that we are not clipping the POY beam. After all that, we will level the table.

ELOG V3.1.3-